OpenStudy (anonymous):

Determine the equation of the tangent line at the value x = 1 for the function f(x) = -5 x + \frac{10}{x} + 4. The slope-intercept form of the equation of the tangent line at the value x = 1 is y = .

OpenStudy (anonymous):

\[ f(x) = -5 x + \frac{10}{x} + 4.\]

OpenStudy (anonymous):

\[f'(x)|_{x=1}= - 5 - \frac{10}{x^2} \implies f'(x)|_{x=1} = -5 - 10 =-15\] \[f(x)|_{x=1} = -5 + 10 + 4 = 9\] \[y - y_1 = m(x - x_1)\] \[y-9 = -15(x - 1)\]

OpenStudy (anonymous):

\[\mathsf{\large \color{red}{\text{FEEL LIKE A SIR}}}\]

OpenStudy (anonymous):

thanks